### v1

@@ 7.3 Planet detection limits at long periods from the SPIRou wPCA radial velocities
We computed planet detection limits from the SPIRou RVs performing an injection recovery test. The method determines the upper limits by injecting signals in the time series at given periods, calculating the GLS periodograms and determining the maximum amplitude that can still be missed with a given probability. For each period, the upper limit is given by the highest amplitude (at the worst phase) that could be present in the time series without creating a peak above the 1% FAP level in the periodogram. The method consists in injecting for each period a sinusoidal Keplerian signal (i.e. circular orbit) at 12 equi-spaced phases. The amplitude (i.e., the planet signal strength) is then increased until the corresponding peak in the GLS periodogram reaches the 1% FAP level. The 1% FAP confidence level was computed from the levels of the highest peaks of 1000 permuted data sets, and the 1% upper limits’ function was obtained by sorting the power of peaks of all periodograms at each period. We tested periods from 0.8 to 10000 days.
Figure 12 displays the projected mass as a function of the period at which 99% and 75% of the planets injected would be detected in the GLS periodogram of the SPIRou data. The degraded upper limits for periods around 1, 2, 180, 360 days are due to the data sampling and an unavoidable consequence of observing at night from a single location and that SPIRou is only mounted in bright time. Our SPIRou RV data would have recovered with a 99% confidence planets with masses larger than 0.1, 1.0 and 4.7 MJ at a periods of 100, 2560 and 5000 days (0.35, 3 and 5 au) respectively.
Kervella et al. (2022) perform a study aimed at detecting companions of Hipparcos catalog stars based on the proper motion anomaly (PMa) derived from Hipparcos and Gaia measurements. Gl 410 is part of that study, it displays a PMa S/N of 0.13. Kervella et al. (2022) derive companion mass of 0.02 MJ (6.3 M⊕) in a 3 to 5 au orbit. The mass of this companion candidate is significantly below the sensitivity capability of our current SPIRou RV measurements.
@@ 8 TESS photometry analysis
The TESS photometry time-series is shown in Fig. 13. One can clearly observe the modulation of the TESS light-curve by the stellar rotation. However, it is not possible to measure the rotation period with high accuracy, because the time span of the observations (27 days) does not fully cover two complete rotation periods. No transits were identified by the TESS automatic pipeline. To further search for transit events, we applied a procedure customized for Gl 410. After removing the outliers using a 3 clipping procedure, we used the package Wotan (Hippke et al., 2019) to flatten the light curve and to remove the effects of variability due to stellar rotation. The second panel of Fig. 13 displays the resultant detrended light curve, it has a scatter of 403 ppm. Using the transit least squares algorithm of Hippke & Heller (2019), we searched for periodic transit events in the detrended light-curve, in particular at the orbital period of planet b. The algorithm retrieves periodic signals at 3.5 and 7 days with a signal detection efficiency (SDE) of the order of 5, which is not high enough to claim a detection (Fig. 14). Further inspection of those signals reveal they are artefacts. The signals are close to the harmonics of the stellar rotation period and likely are related to residuals of the detrending procedure. No signal is detected at period of 6.02 days, suggesting that Gl 410b is not transiting.
@@ 8.1 Transit detection limits
To determine the detection limits in the TESS data, we conducted an injection recovery test of transit events. Following Cortés-Zuleta et al. (2025), using the stellar parameters of Gl 410 and the semi-mayor axis of 0.053 au, we injected 50 000 transits generated with the software batman (Kreidberg, 2015), on which the orbital inclination and planet radius are drawn randomly from uniform distributions: between 85∘ and 90∘ for the inclination and from 0.1 to 10 for the radius. For each synthetic transit, we computed the box least squares (BLS) periodogram131313https://docs.astropy.org/en/stable/timeseries/bls.html implemented in astropy and measured the power of the signal at a period of 6.02 days. Then, we measured the fraction of transit events that are detectable with the BLS periodogram (i.e., power 1000). The results of the injection test are shown in Fig. 15. We find that planets with radius smaller than 1.8 R⊕ are not detectable in the current TESS data, as well as planets in orbits with inclinations lower than 87 deg.
@@ 9.1 Gl 410b in context
Gl 410b sums up to the growing list of sub-Neptune mass planets detected around low-mass stars ( M⊙). In order to situate the Gl 410 planetary system in context, we will concentrate here in discussing the planet detection statistics of the stars in the 0.5 - 0.6 M⊙ stellar mass bin, the stars with a mass equal to that of Gl 410. This stellar mass range has at the time of writing (01.07.2024) according to the NASA exoplanets catalogue141414https://exoplanetarchive.ipac.caltech.edu 212 planet candidates detected (of 5678 exoplanets). Several of those planets are awaiting further characterization, but already 140 have a reported radius measurement, 83 have a planetary mass measurement and 30 have a mass constraint (i.e., ).
@@
The period is always constrained both in transit and RV planet detections. However, in both methods, detection statistics are strongly positively biased to short periods, as planets in short periods are the easiest to detect. At the time of writing, in the M⊙ stellar mass bin, from 212 planet candidates, there are 161 detections at 50 days, with 80 % of those detections at 15 days. These numbers illustrate the observational biases mentioned. To perform a detailed study correcting for biases, it is out of the scope of this discovery paper, however, we can point out, that planets in a 6 days period, such as Gl 410b, are not unusual in the M⊙ stellar mass bin.
@@
Concerning the planet mass, we are interested in discussing how abundant or scarce are planets in the mass range of Gl 410b (8.41.3 M⊕) in orbits with periods 50 days around stars in the M⊙ stellar mass bin. We show in Fig. 16 the histograms of the planet mass distribution for a days (top panel) and for between 10 and 50 days (bottom panel). The histogram includes both mass measurements and lower-limits. In the sub-sample with days, the most abundant planets are those with masses 10 M⊕ with 40% of the detections, followed by planets with mass between 1/3 and 1 MJ with 23% of the detections. Detection biases make planets with masses 10 M⊕ the hardest to detect; the large fraction of detections of planets in the low mass range, means that they should be intrinsically the most numerous. The histograms show a significant decrease on the frequency of planets for the mass bin between 17 M⊕ (Neptune mass) and 100 M⊕ ( 1/3 MJ). As planets in this mass range are more easily detected by RVs than planets with 10 M⊕, the decrease is real and it is not a detection bias. We find that Gl 410b is located in the bin with the highest probability of detection.
@@
In the context of the sub-sample of planets with periods between 10 and 50 days, we find that: i) the highest fraction of detections is the mass bin between 5 and 10 M⊕; ii) there is a decrease in the number of detections of planets with 5 M⊕; and iii) there are very few detections of giant planets. The smaller fraction of low-mass planets could be an observational bias given the fact that to detect planets with 5 M⊕ around M dwarfs at longer periods it is likely harder because the amplitude becomes smaller and it could be comparable to the activity jitter. The very small fraction of giant planets, in contrast, is not an observational bias, must be real, as giant planets are easier to detect by radial velocities. The tentative planet candidate signal at 18.7 days would thus belong to the category of the most frequent planets detected.
@@
The very fact that Gl 410b is located in the bins with the highest number of detections, reinforces the current view that sub-Neptune mass planets are abundant around M dwarfs. The decrease of the occurence rate with increasing planetary mass, it is a property already highlighted in studies covering the entire M dwarf mass domain e.g., Bonfils et al. (2013), Sabotta et al. (2021), Mignon et al. (2025). Current statistics raise the question of why planets with masses between Neptune and 1/3 MJ are scarce at periods shorter than 50 days around stars whose mass is in the M⊙ mass bin. Whether this could be a consequence of the planet formation process, or a product of the interaction of planetary atmospheres with energetic radiation and stellar wind particles due to the high activity levels of M dwarfs is an open question.
@@
As Gl 410b is not transiting, we don’t have a stringent constraint in their radius. From the sample of planets in M⊙ stellar mass bin, days and masses similar as Gl 410b (7 to 10 M⊙) there is at present only a few planets with measurements of the radius for example TOI-1725b ( R⊕, =9.9 M⊕, Essack et al., 2023), and TOI-2018b ( R⊕, =9.2 M⊕, Dai et al., 2023).
@@ 9.2 Stellar activity and the evolution of Gl 410b
Let us conclude this section briefly discussing the implications of stellar activity in the atmospheric evolution of Gl 410b151515 Pending the confirmation of the planet at 2.99 days with new measurements, we not mention it in this discussion. However, the arguments exposed here for Gl 410b would be applicable to this planet given its shorter period.. As presented in Section 2, Gl 410 is an active star. Activity is the manifestation of the presence of stellar magnetic fields. Magnetic fields heat the chromosphere and coronae and drive stellar winds (e.g. Cranmer & Winebarger, 2019). High-energy radiation from stellar flares and X-ray/EUV coronal emissions, and particle fluxes from winds and coronal mass ejections interact with the magnetospheres and the atmospheres of planets in orbits close to the star driving thermal and non-thermal atmospheric escape processes (for recent reviews see Vidotto, 2022; Kubyshkina, 2024).
Gl 410b is located very close to their host star at 0.05 au. It is therefore expected to have an active interaction with the high-energy radiation and the stellar wind particle fluxes from Gl 410. Using the radius and given in Table 1, we obtain that Gl 410b receive 20.4 Earth’s insulation. It is out of the scope of this discovery paper to perform a detailed study on the conditions on which a planet with 9 Earth masses could retain its atmospheres in case of close proximity with an active M-dwarf such as Gl 410, but it is interesting topic that deserves further study. Furthermore, it would be of extreme interest to investigate the influence that stellar irradiation and activity could have in the chemical composition of the planet’s atmosphere and the potential development of abundance anomalies due to the escape of volatile species (e.g., Louca et al., 2023) in single and multi-planetary systems (e.g., Acuña et al., 2022).
@@ 10 Conclusion
In this paper, we presented a radial-velocity study of Gl 410, a nearby (=12 pc) early M dwarf with a mass of 0.55 M⊙ and an age of 480150 Myr 161616This work, see Appendix A.. We have monitored the star in the near-IR with the spectropolarimeter and velocimeter SPIRou and in the optical with the velocimeter SOPHIE. The SPIRou RVs show a robust periodic signal at =6.0200.004 days. The signal is recovered in the raw LBL data and in the two PCA methods used to correct the LBL RVs for systematics and telluric correction residuals. In our baseline solution, the circular orbit model to the SPIRou wPCA data-set (Table 10), the signal has a semi-amplitude of 4.40.7 m s-1, which indicates a sub-Neptune planet with sin of 8.41.3 Earth masses at a 0.05310.0006 au. The same Keplerian signal is recovered in quasi-simultaneous SOPHIE RV measurements after correction for stellar activity using a GP that has been trained on the measurements of the longitudinal magnetic field () obtained with SPIRou. The independent detection of Gl 410b in the SOPHIE RVs provide further evidence for the Keplerian nature of the periodic signal.
@@
We investigated the TESS archive and searched whether Gl 410b could be a transiting planet. We find no evidence for the transit of Gl 410b. A recovery analysis on the TESS photometry indicates that the transit would have been detected if Gl 410b’s inclination would be higher than 87 deg and its radius larger than 1.8 R⊕. Gl 410b is located very close to their host star, it receives 20.4 Earth’s insulation. It would be of great interest to investigate the influence that stellar irradiation and activity could have in the chemical composition of the planet’s atmosphere and in potential atmospheric escape processes.
@@
We find within the SPIRou wPCA RVs tentative evidence for two additional planetary signals at 2.99 and 18.7 days. The 18.7 day signal (consistent with a sub-Neptune planet) is directly seen in the GLS periodograms and it is recovered in a Radvel two-planet model with broad priors for the period. The 2.99 day signal emerges together with the 18.7 day signal in an and apodize sine periodogram (ASP) analysis which takes into account stellar activity indicators measured with SPIRou and a model for the correlated noise. The planet at 2.99 days could be in 2:1 mean motion resonance with the 6.02 days planet.
@@
The detection of Gl 410b shows that infrared RVs are a powerful tool to detect planets around active stars, objects that are often excluded from searches in optical RVs. Care should be taken however to correct/filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. They enabled us to use the RV measurements acquired when the star is at km s-1 of the center of telluric absorption lines, which is important to detect additional planetary signals. Further monitoring of Gl 410 is necessary.
@@ Appendix A On the age of Gl 410
To investigate the possibility of a younger age for Gl 410, we estimated its age using two alternative publicly available gyrochronology codes: stardate (Angus et al. 2019) and gyro-inter (Bouma et al. 2023). Both codes converge in a younger age estimate for Gl 410: stardate suggests an age of 4932 Myr; gyro-inter indicates an age of 326 Myr. Furthermore, we checked the position of Gl 410 in the color-magnitude diagram (CMD). The right panel of Fig. 17 displays the Gaia DR3 (Gaia Collaboration et al. 2023) CMD for Gl 410, field stars, together with empirical CMD sequences from various young associations (Gagné et al. 2021). Gl 410 is located right in between the 400 and 580 Myr sequences. Interpolating in log space, an age of about 480 Myr is obtained for Gl 410. Note however, that the scatter around these empirical sequences is non negligible, and that at younger ages the age distribution is complex and likely non Gaussian. From the empirical sequences, we find that Gl 410 is slightly younger than Praesepe that is somewhat unlikely to be as young as the Pleiades (120 Myr, although the probability is not 0%), and very unlikely to be as young as the Tucana-Horologium association (45 Myr). Based on these considerations, we propose an age of 480150 Myr for Gl 410. Further dedicated work it is necessary to determine the age of Gl 410 with a higher precision, but this study is beyond the scope of this paper. A possible way to narrow-down the uncertainty, would be to map the distribution of reference age populations at this spectral type, reverse the distribution using Bayes theorem and Gl 410’s rotational period and then obtain a non-gaussian age probability distribution for Gl 410.

### v2

@@ 7.3 Planet detection limits at long periods from the SPIRou wPCA radial velocities
We computed the planet detection limits from the SPIRou RVs by performing an injection recovery test. The method determines the upper limits by injecting signals in the time series at given periods, calculating the GLS periodograms, and determining the maximum amplitude that can still be missed with a given probability. For each period, the upper limit is given by the highest amplitude (at the worst phase) that could be present in the time series without creating a peak above the 1% FAP level in the periodogram. The method consists of injecting for each period a sinusoidal Keplerian signal (i.e., circular orbit) at 12 equi-spaced phases. The amplitude (i.e., the planet signal strength) is then increased until the corresponding peak in the GLS periodogram reaches the 1% FAP level. The 1% FAP confidence level was computed from the levels of the highest peaks of 1000 permuted data sets, and the 1% upper limits was obtained by sorting the power of the peaks of all periodograms at each period. We tested periods from 0.8 to 10000 days.
Figure 12 displays the projected mass as a function of the period at which 99% and 75% of the planets injected would be detected in the GLS periodogram of the SPIRou data. The degraded upper limits for the periods around 1, 2, 180, and 360 days are due to the data sampling and are an unavoidable consequence of observing at night from a single location and from the fact that SPIRou is only mounted in bright time. Our SPIRou RV data can detect with a 99% confidence planets with masses larger than 0.1, 1.0 and 4.7 MJ at a periods of 100, 2560 and 5000 days (0.35, 3 and 5 au), respectively.
Kervella et al. (2022) performed a study aimed at detecting companions of Hipparcos catalog stars based on the proper motion anomaly derived from Hipparcos and Gaia measurements. Gl 410 is part of that study, and it displays a proper motion anomaly S/N of 0.13. Kervella et al. (2022) derived a companion mass of 0.02 MJ (6.3 M⊕) in a 3 to 5 au orbit. The mass of this companion candidate is significantly below the sensitivity capability of our current SPIRou RV measurements.
@@ 8 TESS photometry analysis
The TESS photometry time series is shown in Fig. 13. One can clearly observe the modulation of the TESS light-curve by the stellar rotation. However, it is not possible to measure the rotation period with high accuracy because the time span of the observations (27 days) does not fully cover two complete rotation periods. No transits were identified by the TESS automatic pipeline. To further search for transit events, we applied a procedure customized for Gl 410. After removing the outliers using a 3 clipping procedure, we used the package Wotan (Hippke et al., 2019) to flatten the light curve and to remove the effects of variability due to stellar rotation. The second panel of Fig. 13 displays the resultant detrended light curve, which has a scatter of 403 ppm. Using the transit least squares algorithm of Hippke & Heller (2019), we searched for periodic transit events in the detrended light-curve, in particular at the orbital period of planet b. The algorithm retrieved periodic signals at 3.5 and 7 days with a signal detection efficiency (SDE) on the order of 5, which is not high enough to claim a detection (Fig. 14). Further inspection of those signals revealed thay they are artifacts. The signals are close to the harmonics of the stellar rotation period and are likely related to residuals of the detrending procedure. No signal was detected at period of 6.02 days, suggesting that Gl 410b is not transiting.
@@ 8.1 Transit detection limits
To determine the detection limits in the TESS data, we conducted an injection recovery test of transit events. Following Cortés-Zuleta et al. (2025), using the stellar parameters of Gl 410 and the semimajor axis of 0.053 au, we injected 50 000 transits generated with the software batman (Kreidberg, 2015), for which the orbital inclination and planet radius are drawn randomly from uniform distributions: between 85∘ and 90∘ for the inclination and from 0.1 to 10 for the radius. For each synthetic transit, we computed the box least squares periodogram191919https://docs.astropy.org/en/stable/timeseries/bls.html implemented in astropy and measured the power of the signal at a period of 6.02 days. Then, we measured the fraction of transit events that are detectable with the box least squares periodogram (i.e., power 1000). The results of the injection test are shown in Fig. 15. We fond that planets with a radius smaller than 1.8 R⊕ are not detectable in the current TESS data as well as planets in orbits with inclinations lower than 87 deg.
@@ 9.1 Gl 410b in context
Gl 410b sums up to the growing list of sub-Neptune mass planets detected around low-mass stars ( M⊙). In order to place the Gl 410 planetary system in context, in this section we concentrate on discussing the planet detection statistics of the stars in the 0.50.6 M⊙ stellar mass bin, that is, the stars with a mass equal to that of Gl 410. At the time of writing (1 July 2024), this stellar mass range, according to the NASA exoplanets catalog202020https://exoplanetarchive.ipac.caltech.edu, has 212 planet candidates detected (of 5678 exoplanets). Several of those planets are awaiting further characterization, but 140 already have a reported radius measurement, 83 have a planetary mass measurement and, 30 have a mass constraint (i.e., ).
@@
The period is always determined in transit and in RV planet detections. However, in both methods, detection statistics are strongly positively biased to short periods, as planets in short periods are the easiest to detect. At the time of writing, in the M⊙ stellar mass bin, of the 212 planet candidates, there are 161 detections at 50 days, with 80 % of those detections at 15 days. These numbers illustrate the observational biases mentioned. Perfoming a detailed study correcting for biases is out of the scope of this discovery paper; however, we can point out that planets in a 6-day period, such as Gl 410b, are not unusual in the M⊙ stellar mass bin.
@@
Concerning the planet mass, we are interested in discussing how abundant or scarce planets are in the mass range of Gl 410b (8.41.3 M⊕) in orbits with periods 50 days around stars in the M⊙ stellar mass bin. We show in Fig. 16 the histograms of the planet mass distribution for a days (top panel) and for between 10 and 50 days (bottom panel). The histogram includes both mass measurements and lower limits. In the subsample with days, the most abundant planets are those with masses less than 10 M⊕, with 40% of the detections. These are followed by planets with mass between 1/3 and 1 MJ, with 23% of the detections. Detection biases make planets with masses less than 10 M⊕ the hardest to detect. The large fraction of detections of planets in the low mass range means that they should intrinsically be the most numerous. The histograms show a significant decrease on the frequency of planets for the mass bin between 17 M⊕ (Neptune mass) and 100 M⊕ ( 1/3 MJ). As planets in this mass range are more easily detected by RVs than planets with 10 M⊕, the decrease is real, and it is not a detection bias. We find that Gl 410b is located in the bin with the highest probability of detection.
@@
In the context of the subsample of planets with periods between 10 and 50 days, we find that i) the mass bin between 5 and 10 M⊕ has the highest fraction of detections, ii) there is a decrease in the number of detections of planets with 5 M⊕, and iii) there are very few detections of giant planets. The smaller fraction of low-mass planets could be an observational bias given the fact that detecting planets with 5 M⊕ around M dwarfs at longer periods is likely harder because the amplitude becomes smaller, and it could be comparable to the activity jitter. The very small fraction of giant planets, in contrast, is not an observational bias, and it must be real, as giant planets are easier to detect through RVs. The tentative planet candidate signal at 18.7 days would thus belong to the category of the most frequently detected planets.
@@
The very fact that Gl 410b is located in the bins with the highest number of detections reinforces the current view that sub-Neptune mass planets are abundant around M dwarfs. The decrease of the occurrence rate with increasing planetary mass is a property that has already been highlighted in studies covering the entire M dwarf mass domain (e.g., Bonfils et al. (2013), Sabotta et al. (2021), Mignon et al. (2025)). Current statistics raise the question of why planets with masses between Neptune and 1/3 MJ are scarce at periods shorter than 50 days around stars whose mass is in the M⊙ mass bin. Whether this could be a consequence of the planet formation process, or a product of the interaction of planetary atmospheres with energetic radiation and stellar wind particles due to the high activity levels of M dwarfs is an open question.
@@
Since Gl 410b is not transiting, we do not have a stringent constraint on its radius. Based on the sample of planets in the M⊙ stellar mass bin with days and masses similar to Gl 410b (7 to 10 M⊙), at present there are only a few planets for whom measurements of the radius are available, for example, TOI-1725b ( R⊕, =9.9 M⊕, Essack et al., 2023) and TOI-2018b ( R⊕, =9.2 M⊕, Dai et al., 2023).
@@ 9.2 Stellar activity and the evolution of Gl 410b
Let us conclude this section by briefly discussing the implications of stellar activity in the atmospheric evolution of Gl 410b212121Due to the pending the confirmation of the planet at 2.99 days with new measurements, we not mention it in this discussion. However, the arguments exposed here for Gl 410b would be applicable to this planet given its shorter period.. As presented in Section 2, Gl 410 is an active star. Activity is the manifestation of the presence of stellar magnetic fields. Magnetic fields heat the chromosphere and coronae and drive stellar winds (e.g., Cranmer & Winebarger, 2019). High-energy radiation from stellar flares and X-ray/EUV coronal emissions and particle fluxes from winds and coronal mass ejections interact with the magnetospheres and the atmospheres of planets in orbits close to the star, thus driving thermal and non-thermal atmospheric escape processes (for recent reviews see Vidotto, 2022; Kubyshkina, 2024).
As Gl 410b is located very close to its host star at 0.05 au, it is therefore expected to have an active interaction with the high-energy radiation and the stellar wind particle fluxes from Gl 410. Using the radius and given in Table 6, we obtained that Gl 410b receives 20.4 Earth’s insulation. It is out of the scope of this discovery paper to perform a detailed study on the conditions on which a planet with 9 Earth masses could retain its atmospheres in the case that it is in the proximity of an active M dwarf such as Gl 410, but it is an interesting topic that deserves further study. Furthermore, it would be of extreme interest to investigate the influence that stellar irradiation and activity could have in the chemical composition of the planet’s atmosphere and the potential development of abundance anomalies due to the escape of volatile species (e.g., Louca et al., 2023) in single- and multi-planetary systems (e.g., Acuña et al., 2022).
@@ 10 Conclusion
In this paper, we have presented an RV study of Gl 410, a nearby (=12 pc) early M dwarf with a mass of 0.55 M⊙ and an age of 480150 Myr 222222This work, see Appendix A.. We monitored the star in the near-IR with the spectropolarimeter and velocimeter SPIRou and in the optical with the velocimeter SOPHIE. The SPIRou RVs show a robust periodic signal at =6.0200.004 days. The signal is recovered in the raw LBL data and in the two PCA methods used to correct the LBL RVs for systematics and telluric correction residuals. In our baseline solution, the circular orbit model of the SPIRou wPCA data set (Table 16), the signal has a semi-amplitude of 4.40.7 m s-1, which indicates a sub-Neptune planet with sin of 8.41.3 Earth masses at 0.05310.0006 au. The same Keplerian signal is recovered in quasi-simultaneous SOPHIE RV measurements after correction for stellar activity using a GP that was trained on the measurements of the longitudinal magnetic field () obtained with SPIRou. The independent detection of Gl 410b in the SOPHIE RVs provides further evidence for the Keplerian nature of the periodic signal.
@@
We searched the TESS archive and investigated whether Gl 410b could be a transiting planet. We found no evidence for the transit of Gl 410b. A recovery analysis on the TESS photometry indicates that the transit would have been detected if Gl 410b’s inclination were higher than 87 deg and its radius were larger than 1.8 R⊕. Gl 410b is located very close to its host star, and it receives 20.4 times Earth’s insulation. It would be of great interest to investigate the influence that stellar irradiation and activity could have on the chemical composition of the planet’s atmosphere and in the potential atmospheric escape processes.
@@
We find within the SPIRou wPCA RVs, there is tentative evidence for two additional planetary signals at 2.99 and 18.7 days. The 18.7-day signal (which is consistent with a sub-Neptune planet) is directly seen in the GLS periodograms and it is recovered in a RadVel two-planet model with broad priors for the period. The 2.99-day signal emerges together with the 18.7-day signal in an and apodize sine periodogram analysis, which takes into account stellar activity indicators measured with SPIRou and a model for the correlated noise. The planet at 2.99 days could be in a 2:1 mean motion resonance with the 6.02-day planet and could have a mass of a few Earth masses.
@@
The detection of Gl 410b shows that infrared RVs are a powerful tool for detecting planets around active stars, objects which are often excluded from searches in optical RVs. Care should be taken, however, to correct and filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. This approach enabled us to use the RV measurements acquired when the star is at km s-1 of the center of the telluric absorption lines, which is important to detect additional planetary signals. Further monitoring of Gl 410 is necessary.
@@ Appendix A On the age of Gl 410
To investigate the possibility of a younger age for Gl 410, we estimated its age using two alternative publicly available gyrochronology codes: stardate (Angus et al. 2019) and gyro-inter (Bouma et al. 2023). Both codes converge in a younger age estimate for Gl 410: stardate suggests an age of 4932 Myr; gyro-inter indicates an age of 326 Myr. Furthermore, we checked the position of Gl 410 in the color-magnitude diagram (CMD). The right panel of Fig. 17 displays the Gaia DR3 (Gaia Collaboration et al. 2023) CMD for Gl 410, field stars, together with empirical CMD sequences from various young associations (Gagné et al. 2021). Gl 410 is located between the 400 and 580 Myr sequences. Interpolating in log space, an age of about 480 Myr is obtained for Gl 410. We note, however, that the scatter around these empirical sequences is non negligible, that at younger ages the age distribution is complex and likely non Gaussian, and that the age is anchored on the ages assumed for Group X (Oh et al. 2017) and Coma Ber. From the empirical sequences, we find that Gl 410 is slightly younger than Praesepe that is somewhat unlikely to be as young as the Pleiades (120 Myr, although the probability is not 0%), and very unlikely to be as young as the Tucana-Horologium association (45 Myr). Based on these considerations, we propose an age of 480150 Myr for Gl 410. Further dedicated work is necessary to determine the age of Gl 410 with a higher precision, but this study is beyond the scope of this paper. A possible way to narrow-down the uncertainty, would be to map the distribution of reference age populations at this spectral type, reverse the distribution using Bayes theorem and Gl 410’s rotational period and then obtain a non-gaussian age probability distribution for Gl 410.